OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle 169=x^2+y^2 at the point (-12,5).

OpenStudy (anonymous):

Can you calculate the first derivative of the equation for the circle?

OpenStudy (anonymous):

To be honest, I have no clue what that is lol.

OpenStudy (anonymous):

np. y' = +-x/sqrt(169-x^2)^(1/2)

OpenStudy (anonymous):

You will be using "+" since this a point in the second quadrant, so we just put -12 for x in this derivative equation. That is, to get the slope for the tangent line. BTW are you in calculus?

OpenStudy (anonymous):

No, I'm in Algebra2.

OpenStudy (anonymous):

So, slope is +1/5. Okay, we need to do this by another method.

OpenStudy (anonymous):

Do I need to know the center?

OpenStudy (anonymous):

We can't expect you to use methods not available to you at this point. So, this is a circle a tthe origin, point (0, 0). Yes, you need the center. But now we know it. It is implied by the form of the equation.

OpenStudy (anonymous):

So, the next thing we have to do, after identifying the center which is (0, 0), is get the slope of the line from (0, 0) to (-12, 5). Do you know how to do that?

OpenStudy (anonymous):

Do I use slope formula?

OpenStudy (anonymous):

\[\frac{ y _{1} - y _{2} }{ x _{1} - x _{2} } = m\] where m is the slope. And (x1, y1) and (x2, y2) are 2 points on the line. And we know of 2 points on that line, so you would be able to calculate the slope. Can you do this substitution and show me the slope?

OpenStudy (anonymous):

12/5?

OpenStudy (anonymous):

12/5 will be the slope of the tangent line, which is the second step of this problem. The slope of the line between (0, 0) and (-12, 5) is the negative reciprocal or -5/12. Your teacher must have given you a combined step. Anyway, the tangent line slope is 12/5. And now you can use the point-slope formula.

OpenStudy (anonymous):

The first part of the equation is "y=12/5x" correct?

OpenStudy (anonymous):

Well, I'm not sure what "first part" means here, but I believe you are on the right track, because the slope of the tangent line is 12/5, so you have that part identified correctly, if that is what you mean. I'll write out the point-slope formula. It's similar to the other formula I wrote.

OpenStudy (anonymous):

\[\frac{ y - y _{1} }{ x - x _{1} } = m\] or, putting th denominator over on the right side, we have\[y - y _{1} = m(x - x _{1})\]

OpenStudy (anonymous):

You know m is 12/5 and you have the point (-12, 5) for (x1, y1), so you can now just substitute.

OpenStudy (anonymous):

I got y=12/5x+169/5

OpenStudy (anonymous):

Yes! Very good!

OpenStudy (anonymous):

Thank you!